# Reconstruction of frequency-localized functions from pointwise samples via least squares and deep learning

### v1

@@ 5.1 A preliminary result for the main theorems
$\Big{(}\sum_{k=1}^{K}|b_{k}|^{2}\Big{)}^{1/2}=\Big{\|}\sum_{k=1}^{K}b_{k}\varphi^{k}_{\varepsilon}\|_{L^{2}_{\bf u}([-1,1]^{d})}$ $=\Big{\|}\sum_{k=1}^{K}b_{k}(\varphi^{k}_{\varepsilon}-{\rm R}(\Psi^{k}_{\varepsilon}))\Big{\|}_{L^{2}_{\bf u}([-1,1]^{d})}$
$\leq\sum_{k=1}^{K}|b_{k}|\|\varphi^{k}_{\varepsilon}-{\rm R}(\Psi^{k}_{\varepsilon})\|_{L^{2}_{\bf u}([-1,1]^{d})}$
$\leq\varepsilon\sqrt{K}B(d,n)\Big{(}\sum_{k=1}^{K}|b_{k}|^{2}\Big{)}^{1/2}<\Big{(}\sum_{k=1}^{K}|b_{k}|^{2}\Big{)}^{1/2},$
a contradiction. Since, due to (5.70), every element in $\mathcal{N}_{\Lambda,\mathsf{w},\varepsilon}$ is realized as a linear combination of ${\rm R}(\Psi_{\varepsilon}^{k})$, the problem (3.7) is equivalent to finding $(\vec{z})^{\natural}\in\mathbb{C}^{K}$ such that
$(\vec{z})^{\natural}\in\operatorname*{\arg\!\min}_{\vec{z}\in\mathbb{C}^{K}}\|\tilde{A}\vec{z}-\vec{b}\|_{2}^{2}$
where $\vec{b}=\frac{1}{\sqrt{m}}(f(\vec{y}_{j})+\eta_{j})_{j=1}^{m}$. We look at the smallest singular value of $\tilde{A}$ next. By Weyl’s Lemma [22, Theorem 4.3.1], we have
@@
where $\sigma_{j}$ denotes the $j$th singular value of $A$. Hence, for all unit vector $\vec{z}\in\mathbb{C}^{K}$,
$\|(A-\tilde{A})\vec{z}\|_{2}^{2}$ $=\sum_{j=1}^{m}|\sum_{k=1}^{K}(A_{jk}-\tilde{A}_{jk})z_{k}|^{2}$
$\leq\sum_{j=1}^{m}\bigg{(}\sum_{k=1}^{K}|A_{jk}-\tilde{A}_{jk}||z_{k}|\bigg{)}^{2}$
$\leq\sum_{j=1}^{m}\bigg{(}\sum_{k=1}^{K}\frac{1}{\sqrt{m}}\|\varphi_{\mathsf{w},k}-{\rm R}(\Psi^{k}_{\varepsilon})\|_{L^{\infty}([-1,1]^{d})}|z_{k}|\bigg{)}^{2}$
$=\bigg{(}\sum_{k=1}^{K}\|\varphi_{\mathsf{w},k}-{\rm R}(\Psi^{k}_{\varepsilon})\|_{L^{\infty}([-1,1]^{d})}|z_{k}|\bigg{)}^{2}$
$\leq\sum_{k=1}^{K}\|\varphi_{\mathsf{w},k}-{\rm R}(\Psi^{k}_{\varepsilon})\|_{L^{\infty}([-1,1]^{d})}^{2}\|\vec{z}\|_{2}^{2}$
$\leq KB(d,n)^{2}\varepsilon^{2},$
@@
$\|A-\tilde{A}\|_{2}\leq\sqrt{K}B(d,n)\varepsilon.$ (5.75)
We have demonstrated in the proof of Theorem 3.1 that $\sigma_{\rm min}(A)>\sqrt{1-\delta}$ with probability at least $1-\beta$, as long as $m\geq c_{\delta}\kappa^{\star}\log(\texttt{\#}\Lambda/\beta)$. Hence, from (5.74), (5.75), we get
$\sigma_{\rm min}(\tilde{A})\geq\sigma_{\rm min}(A)-\|A-\tilde{A}\|_{2}\geq\sigma_{\rm min}(A)-\sqrt{K}B(d,n)\varepsilon>\sqrt{1-\delta}-\sqrt{K}B(d,n)\varepsilon,$ (5.76)
with the same probability. Thus, by choosing $\varepsilon\in(0,1/(\sqrt{K}B(d,n)))$ additionally sufficiently small such that
$\sqrt{K}B(d,n)\varepsilon\leq\frac{1}{2}\sqrt{1-\delta},$
@@
By applying the least squares approximation theory from Subsection 2.2, specifically leveraging (5.2) and the lower bound given in (5.77), we conclude that (3.7) has, with probability at least $1-\beta$, a unique solution $\Psi^{\natural}$ such that
@@
$\leq\|f-g\|_{L^{\infty}([-1,1]^{d})}+\sqrt{K}B(d,n)\|g\|_{L^{2}_{\bf u}([-1,1]^{d})}\varepsilon,$
@@ 6 Conclusion
We conclude by discussing some limitations of our work and indicating possible directions for future research. First, our theoretical results hold in dimensions $d=1,2,3$. In principle, our arguments can be applied to higher dimensions, but the results would be affected by the curse of dimensionality. This arises due to the following reasons. The sample complexity bound in $m$ of Theorem 3.1 relies on the estimation of an upper bound on $\kappa(S_{\Lambda})$, which is given in Proposition 5.8. This upper bound estimate stems from Proposition 5.9 and Lemma 5.10, both providing dimension-dependent bounds for the cardinality of the hyperbolic cross index set and its slices. Specifically, (5.6) holds only beyond a certain threshold $n_{\star}(d)$. For $d\geq 3$, $n_{\star}(d)$ is no longer guaranteed to be one. Additionally, (5.23) applies after another distinct threshold in $n$ for each $d$. Since both (5.6) and (5.23) do not necessarily hold for all $n\in\mathbb{N}$ for each $d$, this disrupts our inductive reasoning across dimensions, which is intricately tied to the geometry of hyperbolic crosses.
@@
While Proposition 5.9 and Lemma 5.10 are integral to our proof strategy for circumventing the dimensional impact in the upper bound of $\kappa(\mathcal{S}_{\Lambda})$, the potential for the curse of dimensionality to emerge remains evident, as $\kappa(\mathcal{S}_{\Lambda})\geq\|\varphi_{\mathsf{w},\vec{0}}\|_{L^{\infty}([-1,1]^{d})}^{2}$, and the latter quantity is of order $\mathcal{O}(\mathsf{w}^{d})$, by Proposition 5.2. A potential fix to mitigate this problem is to consider a preconditioning scheme (see [3, Chapter 5]) where samples are drawn from an orthogonal measure on $[-1,1]^{d}$, alongside leveraging hyperbolic crosses of anisotropic type. Alternatively, recovery strategies based on compressed sensing (see [3, Chapter 7] and the references therein) could be explored. These considerations fall outside the scope of this paper, but they present promising directions for future research.
@@
The error bounds in Theorems 3.1 and 3.2 depend on the best approximation error of the target function $f$ with respect to the Slepian basis $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$. It is possible to obtain approximation rates that depend on the number of samples $m$ for specific function classes, similar to the approach in [2] for analytic functions. To address this approximation-theoretical challenge, one would need to define a suitable class of (approximately) bandlimited functions for which the best approximation error $\inf_{g\in\mathcal{S}_{\Lambda}}\|f-g\|_{L^{\infty}([-1,1]^{d})}$ can be explicitly estimated in terms of #$\texttt{\#}\Lambda$. This type of bound is not readily available in the literature and so would need to be derived.
@@
Additionally, as noted in Section 3, the NNs in the class $\mathcal{N}_{\Lambda,\mathsf{w},\varepsilon}$ of Theorem 3.2 are not fully trained like those considered in our numerical experiments. In fact, their first-to-second-last layers are explicitly constructed to emulate the Slepian basis, and the last layer is trained via least squares. Addressing this gap between theory and practice is an important direction for future work. One possible approach is given by the proof strategy of [5, Theorem 3.2], which shows that training problems based on any family of fully connected NNs possess uncountably many minimizers that achieve accurate approximations.
@@
Finally, our numerical experiments in dimensions one and two demonstrate that both reconstruction methods using least squares and deep learning improve as the number of training samples increases. Although least squares approximation is far more effective in dimension one, deep learning can match and slightly surpass it in dimension two. However, identifying an NN architecture that consistently achieves superior performance across different benchmark functions is an open question that deserves a more comprehensive numerical investigation.
@@ Acknowledgments
AMN is supported by the Austrian Science Fund (FWF) Project P-37010. SB and JJB were partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through grants RGPIN-2020-06766 and RGPIN-2023-04244, respectively, and the Fonds de Recherche du Québec Nature et Technologies (FRQNT) through grants 313276 and 340894, respectively.
@@ Appendix A Basic neural network calculus
Proposition A.1 ([30, Proposition 2.5]).
Let $\varepsilon\in(0,1)$, $B>0$, and ×$\times:\mathbb{R}^{2}\to\mathbb{R}$ be defined by ×$\times(x,y)=xy$. Then, there exists a ReLU NN ×~$\widetilde{\times}_{\varepsilon,B}$ such that ${\rm R}(\widetilde{\times}_{\varepsilon,B}):[-B,B]^{2}\to\mathbb{R}$ and that
@@
$L(\widetilde{\times}_{\varepsilon,B})\leq C(1+\log_{2}(B/\varepsilon))\quad\text{ and }\quad W(\widetilde{\times}_{\varepsilon,B})\leq C(1+\log_{2}(B/\varepsilon)).$
@@
Definition A.2 ([34, Definition 2.2]).
@@
Definition A.3 ([34, Definition 2.7]).
@@
Proposition A.6 ([35, Lemma 5.4]).
@@
Our final result addresses the polynomial approximability of NNs and is derived from [30, Proposition 2.11].
@@
and for $k=0$, an NN $\Phi^{0}_{\varepsilon}$ satisfying $P_{0}={\rm R}(\Phi^{0}_{\varepsilon})=1$ on $[-1,1]$. Moreover, there exists a universal constant $C>0$ such that for $k\in\mathbb{N}$

### v2

@@ 5.1 A preliminary result for the main theorems
$\Big(\sum_{\vec{\nu}\in\Lambda}|b_{\vec{\nu}}|^{2}\Big)^{1/2}=\Big\|\sum_{\vec{\nu}\in\Lambda}b_{\vec{\nu}}\varphi^{\vec{\nu}}_{\varepsilon}\|_{L^{2}_{\bf u}([-1,1]^{d})}$ $=\Big\|\sum_{\vec{\nu}\in\Lambda}b_{\vec{\nu}}(\varphi^{\vec{\nu}}_{\varepsilon}-{\rm R}(\Psi^{\vec{\nu}}_{\varepsilon}))\Big\|_{L^{2}_{\bf u}([-1,1]^{d})}$
$\leq\sum_{\vec{\nu}\in\Lambda}|b_{\vec{\nu}}|\|\varphi^{\vec{\nu}}_{\varepsilon}-{\rm R}(\Psi^{\vec{\nu}}_{\varepsilon})\|_{L^{2}_{\bf u}([-1,1]^{d})}$
$\leq\varepsilon\sqrt{\texttt{\#}\Lambda}\,B(d,n)\Big(\sum_{\vec{\nu}\in\Lambda}|b_{\vec{\nu}}|^{2}\Big)^{1/2}<\Big(\sum_{\vec{\nu}\in\Lambda}|b_{\vec{\nu}}|^{2}\Big)^{1/2},$
a contradiction. Since, due to (5.67), every element in $\mathcal{N}_{\Lambda,\mathsf{w},\varepsilon}$ is realized as a linear combination of ${\rm R}(\Psi_{\varepsilon}^{\vec{\nu}})$, the problem (3.7) is equivalent to finding $(\vec{z})^{\natural}\in\mathbb{C}^{\texttt{\#}\Lambda}$ such that
$(\vec{z})^{\natural}\in\operatorname*{\arg\!\min}_{\vec{z}\in\mathbb{C}^{\texttt{\#}\Lambda}}\|\tilde{A}\vec{z}-\vec{b}\|_{2}^{2}$
where $\vec{b}=\frac{1}{\sqrt{m}}(f(\vec{y}_{j})+\eta_{j})_{j=1}^{m}$. We look at the smallest singular value of $\tilde{A}$ next. By Weyl’s Lemma [horn2012matrix, Theorem 4.3.1], we have
@@
where $\sigma_{j}$ denotes the $j$th singular value of $A$. Hence, for all unit vector $\vec{z}\in\mathbb{C}^{\texttt{\#}\Lambda}$,
$\|(A-\tilde{A})\vec{z}\|_{2}^{2}$ $=\sum_{j=1}^{m}|\sum_{\vec{\nu}\in\Lambda}(A_{j,\vec{\nu}}-\tilde{A}_{j,\vec{\nu}})z_{\vec{\nu}}|^{2}$
$\leq\sum_{j=1}^{m}\bigg(\sum_{\vec{\nu}\in\Lambda}|A_{j,\vec{\nu}}-\tilde{A}_{j,\vec{\nu}}||z_{\vec{\nu}}|\bigg)^{2}$
$\leq\sum_{j=1}^{m}\bigg(\sum_{\vec{\nu}\in\Lambda}\frac{1}{\sqrt{m}}\|\varphi_{\mathsf{w},\vec{\nu}}-{\rm R}(\Psi^{\vec{\nu}}_{\varepsilon})\|_{L^{\infty}([-1,1]^{d})}|z_{\vec{\nu}}|\bigg)^{2}$
$=\bigg(\sum_{\vec{\nu}\in\Lambda}\|\varphi_{\mathsf{w},\vec{\nu}}-{\rm R}(\Psi^{\vec{\nu}}_{\varepsilon})\|_{L^{\infty}([-1,1]^{d})}|z_{\vec{\nu}}|\bigg)^{2}$
$\leq\sum_{\vec{\nu}\in\Lambda}\|\varphi_{\mathsf{w},\vec{\nu}}-{\rm R}(\Psi^{\vec{\nu}}_{\varepsilon})\|_{L^{\infty}([-1,1]^{d})}^{2}\|\vec{z}\|_{2}^{2}$
$\leq\texttt{\#}\Lambda\,B(d,n)^{2}\varepsilon^{2},$
@@
$\|A-\tilde{A}\|_{2}\leq\sqrt{\texttt{\#}\Lambda}\,B(d,n)\varepsilon.$ (5.72)
We have demonstrated in the proof of Theorem 3.1 that $\sigma_{\rm min}(A)>\sqrt{1-\delta}$ with probability at least $1-\beta$, as long as
$m\geq((1-\delta)\log(1-\delta)+\delta)^{-1}(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}\log(\texttt{\#}\Lambda/\beta).$
Hence, from (5.71), (5.72), we get
$\sigma_{\rm min}(\tilde{A})\geq\sigma_{\rm min}(A)-\|A-\tilde{A}\|_{2}\geq\sigma_{\rm min}(A)-\sqrt{\texttt{\#}\Lambda}\,B(d,n)\varepsilon>\sqrt{1-\delta}-\sqrt{\texttt{\#}\Lambda}\,B(d,n)\varepsilon,$ (5.73)
with the same probability. Thus, by choosing $\varepsilon\in(0,1/(\sqrt{\texttt{\#}\Lambda}\,B(d,n)))$ additionally sufficiently small such that
$\sqrt{\texttt{\#}\Lambda}\,B(d,n)\varepsilon\leq\frac{1}{2}\sqrt{1-\delta},$
@@
By applying the least squares approximation theory from Section 2.2 and Section 5.3, specifically leveraging (5.15), (5.3) and the lower bound given in (5.74), we conclude that (3.7) has, with probability at least $1-\beta$, a unique solution $\Psi^{\natural}$ such that
@@
$\leq\|f-g\|_{L^{\infty}([-1,1]^{d})}+\sqrt{\texttt{\#}\Lambda}\,B(d,n)\|g\|_{L^{2}_{\bf u}([-1,1]^{d})}\varepsilon,$
@@ 6 Conclusion
We conclude by discussing some limitations of our work and indicating possible directions for future research. First, our theoretical results hold in dimensions $d=1,2,3$. In principle, our arguments can be applied to higher dimensions, but the results would be affected by the curse of dimensionality. This arises due to the following reasons. The sample complexity bound in $m$ of Theorem 3.1 relies on the estimation of an upper bound on $\kappa(S_{\Lambda})$, which is given in Proposition 5.8. This upper bound estimate follows from Proposition 5.9, whose proof provides implicit dimension-dependent bounds for the cardinalities of the slices of the hyperbolic cross index set. An earlier version of the proof of Proposition 5.9 relied on the following estimate for the cardinality of the hyperbolic cross, for which an asymptotic version can be found in [dobrovol1998number].
Lemma 6.1 (See [chernov2016new, Theorem 3.5] and [adcock2022sparse, Lemma B.3]).
$\texttt{\#}\Lambda^{\rm HC}_{n-1}<\frac{n(\log n+d\log 2)^{d-1}}{(d-1)!}.$
An estimate for $n_{\star}(d)$ can be derived, for example, from the calculations presented in the proof of [chernov2016new, Theorem 2.4]. In particular, we can take $n_{\star}(2)=1$. However, for $d\geq 3$, $n_{\star}(d)$ is no longer guaranteed to be $1$. Moreover, (5.19) may only hold above a separate threshold in $n$ for each $d$. This prevents a straightforward inductive argument across dimensions that involves the intricate geometry of the hyperbolic cross in higher dimensions.
@@
While Proposition 5.9 is integral to our proof strategy for circumventing the dimensional impact in the upper bound of $\kappa(\mathcal{S}_{\Lambda})$, the potential for the curse of dimensionality to emerge remains evident, as $\kappa(\mathcal{S}_{\Lambda})\geq\|\varphi_{\mathsf{w},\vec{0}}\|_{L^{\infty}([-1,1]^{d})}^{2}$, and the latter quantity is of order $\mathcal{O}(\mathsf{w}^{d})$, by Proposition 5.2. A potential fix to mitigate this problem is to consider a preconditioning scheme (see [adcock2022sparse, Chapter 5]) where samples are drawn from an orthogonal measure on $1,1]^{d}$, alongside leveraging hyperbolic crosses of anisotropic type. Alternatively, recovery strategies based on compressed sensing (see [adcock2022sparse, Chapter 7] and the references therein) could be explored. These considerations fall outside the scope of this paper, but they present promising directions for future research.
@@
The error bounds in Theorems 3.1 and 3.3 depend on the best approximation error of the target function $f$ with respect to the Slepian basis $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$. It is possible to obtain approximation rates that depend on the number of samples $m$ for specific function classes, similar to the approach in [adcock2022deep] for analytic functions. To address this approximation-theoretical challenge, one would need to define a suitable class of (approximately) bandlimited functions for which the best approximation error $\inf_{g\in\mathcal{S}_{\Lambda}}\|f-g\|_{L^{\infty}([-1,1]^{d})}$ can be explicitly estimated in terms of $\texttt{\#}\Lambda$. This type of bound is not readily available in the literature and so would need to be derived. Another interesting open question is whether the error bound in (3.3) holds when the inf is taken over the neural networks in $\mathcal{N}_{\Lambda,\mathsf{w},\varepsilon}$ as opposed to $\mathcal{S}_{\Lambda}$.
@@
Additionally, as noted in Section 3, the NNs in the class $\mathcal{N}_{\Lambda,\mathsf{w},\varepsilon}$ of Theorem 3.3 are not fully trained like those considered in our numerical experiments. In fact, their first-to-second-last layers are explicitly constructed to emulate the Slepian basis, and the last layer is trained via least squares. Addressing this gap between theory and practice is an important direction for future work. One possible approach is given by the proof strategy of [adcock2024optimal, Theorem 3.2], which shows that training problems based on any family of fully connected NNs possess uncountably many minimizers that achieve accurate approximations.
@@
Finally, our numerical experiments in dimensions one and two demonstrate that both reconstruction methods using least squares and deep learning improve as the number of training samples increases. Although least squares approximation is far more effective in dimension one, deep learning can match and slightly surpass it in dimension two. However, identifying an NN architecture that consistently achieves superior performance across different benchmark functions is an open question that deserves a more comprehensive numerical investigation. Additionally, our experiments in Section 4.2 show that a novel Slepian-based initialization—based on the construction of Theorem 3.3—can outperform standard random initializations when approximating a one-dimensional frequency localized test function. Carrying out a more extensive experimentation to see if this advantage can be seen for more functions and in higher dimensions is an interesting avenue of future work.
@@ Acknowledgments
AMN is supported by the Austrian Science Fund (FWF) Project P-37010. SB and JJB were partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through grants RGPIN-2020-06766 and RGPIN-2023-04244, respectively, and the Fonds de Recherche du Québec Nature et Technologies (FRQNT) through grants 313276, 340894 and 359708, respectively.
@@ Appendix A Basic neural network calculus
Proposition A.1 ([opschoor2022exponential, Proposition 2.5]).
Let $\varepsilon\in(0,1)$, $B>0$, and $\times:\mathbb{R}^{2}\to\mathbb{R}$ be defined by $\times(x,y)=xy$. Then, there exists a ReLU NN $\widetilde{\times}_{\varepsilon,B}$ such that ${\rm R}(\widetilde{\times}_{\varepsilon,B}):[-B,B]^{2}\to\mathbb{R}$ and that
@@
[…]
@@
Definition A.2 ([PetV2018OptApproxReLU, Definition 2.2]).
@@
Definition A.3 ([PetV2018OptApproxReLU, Definition 2.7]).
@@
Proposition A.6 ([petersen2024mathematical, Lemma 5.4]).
@@
Our final result addresses the polynomial approximability of NNs and is derived from [opschoor2022exponential, Proposition 2.11].
@@
and for $k=0$, an NN $\Phi^{0}_{\varepsilon}$ satisfying $P_{0}={\rm R}(\Phi^{0}_{\varepsilon})=1$ on [1,1][-1,1]. Moreover, there exists a universal constant $C>0$ such that for $k\in\mathbb{N}$